FTMW SPECTROSCOPY OF SILYL MERCAPTAN, H_{3}SiSH

\underline{S. THORWIRTH}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; V. LAT-TANZI, OSCAR MARTINEZ, JR., MICHAEL C. MCCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and School of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A.; LI-HONG XU, Department of Physics, Centre for Laser, Atomic and Molecular Studies (CLAMS) University of New Brunswick, Saint John, New Brunswick, Canada E2L 4L5.

By means of Fourier transform microwave spectroscopy of a supersonic jet, the pure rotational spectrum of silyl mercaptan, H_{3}SiSH, has been observed for the first time in a low-current dc discharge through a mixture of silane and hydrogen sulfide heavily diluted in neon. The spectroscopic identification was based on predictions from high-level quantum-chemical calculations at the CCSD(T) level of theory in combination with large basis sets performed using the CFOUR suite of programs^{a}. In addition to the parent isotopic species, rotation lines of the rare isotopologs H_{3}^{29}SiSH, H_{3}^{30}SiSH, H_{3}Si^{34}SH, and D_{3}SiSD have also been observed.

aCFOUR, Coupled-Cluster techniques for Computational Chemistry, a quantum-chemical program package by J.F. Stanton, J.

Gauss, M.E. Harding, P.G. Szalay et al., for the current version, see http://www.cfour.de.